OpenStudy (anonymous):

Here's an easy one for you! I keep coming up with -10=x, but I know that's not the answer. Can anyone please give me a hand? (: 2(x - 4) - 12 = 2 - 3(x + 4)

OpenStudy (anonymous):

2x-8-12=2-3x-12 2x-8=2-3x -8-2=-3x-2x -10=-5x x=2

OpenStudy (anonymous):

Thank you! That makes sense!

OpenStudy (anonymous):

:)